OpenStudy (anonymous):

I understood none of today's lesson and I need an explanation on a couple problems, the problems where you have to take an equation with multiple variables and make it so one side has a single isolated variable. Please explain these: 2m - nx = x + 4 solve for x x/a - 1 = y/b solve for x V=1/3 (pi) r^2 (h) solve for H A= (f+g/2) h solve for g I only need explanations.

OpenStudy (jchick):

2m - nx = x + 4

OpenStudy (jchick):

2m = x + 4 + nx

OpenStudy (jchick):

2m - 4 = x + nx

OpenStudy (jchick):

2m - 4 = x(1 + n)

OpenStudy (jchick):